# InvFussion: Bridging Supervised and Zero-shot Diffusion for Inverse Problems

### v1

@@ 3 Method
DDRM [31] $22.49$ $23.18$ $23.16$ $12.62$ $53.55$ $53.55$ $10.76$ $72.74$ $72.74$ $25$
DDNM [55] $23.98$ $42.52$ $42.34$ $16.18$ $54.59$ $54.59$ $11.93$ $131.37$ $131.37$ $100$
DPS [9] $18.31$ $9.66$ $17.73$ $10.06$ $16.96$ $67.46$ $9.57$ $26.86$ $35.93$ $1000$*
$\Pi$GDM [47] $22.12$ $6.29$ $6.32$ $15.20$ $12.81$ $12.81$ $10.71$ $19.13$ $19.13$ $100$*
DAPS [56] $17.50$ $143.45$ $60.33$ $16.31$ $16.91$ $19.12$ $11.95$ $56.30$ $87.18$ $1000$
@@
Palette [42] $20.97$ $4.70$ $5.83$ $16.97$ $5.32$ $5.45$ $12.50$ 6.63 $25.54$ $63$
InvFussion $22.59$ 4.65 4.69 $17.14$ 5.08 5.11 $14.89$ $8.33$ 8.41 $63$
MMSE estimation methods (Training-based)
Palette [42] $23.47$ $19.65$ $15.51$ $1$
InvFussion 24.74 - - 19.81 - - 17.62 - - $1$
@@
DDRM [31] $21.21$ $26.12$ $26.22$ $12.21$ $30.49$ $30.49$ $10.60$ $32.89$ $32.89$ $25$
DDNM [55] $22.45$ $69.47$ $69.27$ $15.08$ $95.27$ $95.27$ $11.32$ $115.32$ $115.32$ $100$
DPS [9] $18.26$ $8.96$ $12.86$ $10.36$ $15.23$ $17.74$ $9.52$ $12.80$ $12.80$ $1000$*
$\Pi$GDM [47] $19.99$ $7.12$ $7.14$ $14.29$ $8.74$ $8.74$ $10.59$ $8.93$ $8.93$ $100$*
DAPS [56] $17.32$ $125.73$ $62.81$ $15.35$ $16.31$ $19.66$ $11.65$ $28.27$ $37.03$ $1000$
@@
Palette [42] $19.61$ $6.60$ $6.97$ $15.56$ 6.20 6.20 $12.34$ 6.70 $11.44$ $63$
InvFussion $20.80$ 6.21 6.21 $15.58$ $6.44$ $6.44$ $14.38$ $6.87$ 6.87 $63$
MMSE estimation methods (Training-based)
Palette [42] $22.25$ $18.32$ $15.28$ $1$
InvFussion 23.19 - - 18.34 - - 17.12 - - $1$
    Methods that use the derivative of the network, leading to more computation and memory per NFE.
3.1 MMSE Estimator
Our InvFussion architecture can be used for more than posterior sampling. Having an architecture that make use of the degradation ${\bm{H}}$ enables training a single model with
to produce MMSE predictors for many different degradations. Alternatively, the MMSE predictor can be approximated by a diffusion model trained with Eq. 8 by applying a single denoising step on white Gaussian noise. That is, for sufficiently small $\bar{\alpha(t)}$ we have that ${\mathbf{x}}_{t}\approx{\mathbf{\epsilon}}$, which is independent of ${\mathbf{x}}_{0}$, so that
which also holds for the approximate models being trained. In other words, any conditional diffusion model can be used as an efficient MMSE predictor. We find that this works well (see App. B), and utilize this in our experiments.
@@ 4 Experiments
We evaluate the advantage of InvFussion over existing inverse problem solvers on $64\times 64$ images from the FFHQ [27] and ImageNet [13] datasets. In these experiments, we conduct a comparative analysis of InvFussion against alternative methods that can operate on several types of degradations with a single model. This includes zero-shot methods, utilizing an unconditional model which we trained with an equivalent architectural configuration. Additionally, we evaluate our approach against a “Palette”-style methodology [42] that incorporates the ${\bm{H}}^{\dagger}{\mathbf{y}}$ as input without additional degradation information. We train this model ourselves using the same underlying architecture, to enable a comprehensive assessment of our method’s capabilities disentangled from architectural considerations. The models trained on ImageNet are also class-conditional, and use CFG [25] with a factor of $2.0$ for sampling.
@@
We train and evaluate our model on several categories of inverse problems; strided motion blur (combining various degrees of super-resolution and de-blurring), in-painting by leaving only patches of various sizes, and a general matrix degradation operator. The parameters of each degradation are randomized, to create a large corpus of possible degradations (as detailed in App. A). All conditional models are also trained on unconditional generation by using a degradation which always outputs zeros. Figure 3 shows qualitative comparisons of different inverse problem solving methods for the two datasets. All training-based approaches use the deterministic sampler from EDM [28] along with identical seeds, highlighting the effects of different models.
@@
For quantitative comparisons, we use PSNR to measure the image fidelity, i.e. how close the restored output is to the original. To measure image realism, we use 10K FID [24]. These two metrics are at odds according to the Perception-Distortion-Tradeoff [6], and different methods excel at either of the metrics. We also find that some restoration algorithms sacrifice consistency with the measurements ${\mathbf{y}}$ to accommodate better image realism, in effect ignoring the inverse problem. To measure the image realism of valid solutions to the given inverse problem, we offer to measure the FID of generated samples that have been projected to be consistent with the measurements ${\mathbf{y}}$, which we refer to as C-FID. In this way, we penalize models for the distance between the generated sample and the nearest valid solution to the formulated problem. Tables 1 and 2 show our qualitative analysis on the FFHQ [27] and ImageNet [13] datasets respectively. Our model achieves the best C-FID among all training-based and zero-shot methods in all but a single case, suggesting it is SOTA in generating samples from the posterior. Our InvFussion MMSE estimator based on the same model achieves the best PSNR of all models, while the InvFussion posterior sampler is also among the highest in PSNR. This is also reflected in the qualitative examples for MMSE estimation in  Fig. 4.
Figure 4: Examples comparing MMSE estimators and samples from the posterior for InvFussion against a baseline that is not degradation-aware. InvFussion model is equivalent or superior in both posterior sampling and MMSE estimation.
@@
Interestingly, the gap in performance between InvFussion and Palette highly varies between different degradation families. In problems like in-painting, the degradation is easily inferrable by the network, as the ${\bm{H}}^{\dagger}{\mathbf{y}}$ input to the network contains masked regions (see Fig. 3). Thus in in-painting, we see marginal to no advantage for InvFussion over Palette. On the other hand, in motion-blur and even more so on general matrix degradations, the Palette model struggles to infer the exact degradation, leading to a drop in both PSNR and C-FID, sometimes, even lower than some zero-shot methods which are degradation-aware. This supports our hypothesis that knowledge of the degradation operator, whether inferred or explicit, is critical for correct restoration.
@@
To demonstrate InvFusion’s scalability and performance at higher resolutions, we conducted additional experiments using $256\times 256$ images from FFHQ. The complete results are presented in App. C, with representative examples shown in Fig. 1.
@@
Table 3: Comparison of restoration on a degradation that did not appear in training. InvFussion demonstrates strong adaptation capabilities through its degradation-aware architecture.
@@
DDRM [31] $11.59$ $24.85$ $24.85$
DDNM [55] $12.69$ $75.81$ $75.81$
DPS [9] $10.09$ $8.78$ $14.10$
$\Pi$GDM [47] $11.28$ $7.26$ $7.26$
DAPS [56] $12.12$ $40.84$ $39.53$
@@
Palette [42] $10.97$ $19.91$ $19.89$
InvFussion $11.85$ 5.73 5.73
MSE estimation methods (Training-based)
Palette [42] $13.58$
InvFussion 14.71 - -
Table 4: Comparison of unconditional sampling.
@@
[…]
Baseline [11] 6.66 8.98
Palette [42] 6.56 8.26
InvFussion 6.27 8.19
@@ 4.1 Out-Of-Distribution Degradations and Unconditional Sampling
To measure the generalization capabilities of different methods we evaluate performance on a degradation that is slightly outside of the training distribution. In this case, we perform out-painting from a single rectangular area in the image. The results in Tab. 3, show that Palette completely fails to generalize, despite the similarity to the in-painting task the model has been trained on and the ease at which the degradation operator can be inferred from the input. On the other hand, InvFussion does quite well on this problem, outperforming all zero-shot methods, which are not penalized by the out-of-distribution (OOD) degradation, as they are not trained on any degradation. Finally, we also evaluate all trained models for unconditional generation (on which they had been trained, by using ${\bm{H}}\in R^{0\times D}$). Surprisingly, we find that both inverse-problem models outperform the unconditional baseline, with InvFussion leading with a $7\%$ improvement over the baseline (unconditional diffusion). This is despite the conditional models having no additional capacity, as both the additional model inputs and the Feature Degradation Layer are always output zeros for unconditional sampling. We hypothesize that being exposed to partial image information during training helps the diffusion model converge faster.
@@ 4.2 Noisy Degradations
Following the noiseless linear case explored in 4, we also conduct experiments on noisy inverse problems. We retrain the FFHQ64 model, incorporating additive white Gaussian noise with a standard deviation of $\sigma_{n}=0.1$ in the measurements ${\mathbf{y}}$. The InvFussion architecture remains unchanged, using only the transforms ${\bm{H}}$ and ${\bm{H}}^{\dagger}$ of the degradation – as adding randomly sampled noise at each layer could lead to accumulation of noise. Projection operators are not well-defined for noisy degradations, so only PSNR and FID are evaluated for the noisy case. Our findings in Tab. 8 (App. C) demonstrate the advantage of InvFussion over alternative methods in the noisy settings.
@@ 5 Neural Posterior Principal Components
Beyond the improvements in restoration fidelity and realism, using an operator-aware architecture unlocks several capabilities that are useful for down stream applications. A notable example, is the computation of Neural Posterior Principal Components (NPPCs) [38], which are efficient estimations of the posterior distribution’s largest principal vectors. These direction are meaningful for uncertainty quantification among many other tasks [36, 18], and may have high implications for medical analysis [5, 17]. In the original work by Nehme et al. [38], a network is trained to approximate the posterior principal components for a single pre-defined degradation. Using InvFussion, we can use the same approach to scale up NPPC training to multiple degradations at once, enabling flexible adaptation to the degradation at runtime. We experiment with applying the exact same NPPC loss from [38], using the InvFussion architecture to condition the network on the degradation operator ${\bm{H}}$ for the in-painting of a single area of the image. The examples in Figure 5 shows that InvFussion meaningfully extends NPPC computation for multiple degradation, making a single model viable for a wide range of problems.
@@ 6 Discussion, Limitations and Conclusion
Despite achieving state-of-the-art performance in posterior sampling and MMSE estimation, InvFussion faces several important limitations. A primary constraint lies in the model’s training scope – although InvFussion demonstrates remarkable adaptability to various degradations, it can only be trained on a finite set of degradation types. Even in the linear case, the space of possible degradations scales quadratically with image dimensions, creating fundamental constraints that would be impractical to address solely through increased model capacity and training time. Additionally, while our current results focus exclusively on linear degradations, we anticipate that InvFussion could potentially generalize to non-linear degradations when appropriately trained, though this remains to be empirically validated. A final consideration is the computational and memory intensity of repeated degradation applications, particularly during the training phase. Although this computational overhead may be relatively minor compared to certain zero-shot methods that require additional steps or model derivatives, optimizing these resource requirements represents a key area for future improvement.
Figure 5: Examples of Using InvFussion for NPPC. By being degradation-aware, the model can be trained to predict the MMSE and several leading principle components ${\mathbf{w}}_{i}$ (shown on the left, contrast enhanced) for many degradations.
This work introduces InvFussion, a novel architectural framework that bridges the long-standing gap between training-based and zero-shot approaches in solving inverse problems with diffusion models. By incorporating the degradation operator directly into the network architecture through attention-based layer unrolling, our method achieves state-of-the-art performance while maintaining the flexibility to handle diverse degradation scenarios. The empirical results demonstrate superior performance on multiple dataset across various inverse problems, while offering computational efficiency comparable to existing methods. Beyond its primary application, InvFussion’s capability as a general MMSE estimator and its potential for NPPC estimation opens new avenues for downstream applications. While certain limitations remain, this work represents a significant step forward in the field of accurate and efficient inverse problem solving.
Acknowledgments

### v2

@@ 3.1 MMSE Estimator
DDRM [kawar2022denoising] $22.5$ $23.2$ $23.2$ $12.6$ $53.6$ $53.6$ $10.8$ $72.7$ $72.7$ $25$
DDNM [wang2022zero] $24.0$ $42.5$ $42.3$ $16.2$ $54.6$ $54.6$ $11.9$ $131.4$ $131.4$ $100$
DPS [chung2023diffusion] $18.3$ $9.7$ $17.7$ $10.1$ $17.0$ $67.5$ $9.6$ $26.7$ $35.9$ 10001000*
[…]
DAPS [zhang2024improving] $17.5$ $143.5$ $60.3$ $16.3$ $16.9$ $19.1$ $12.0$ $56.3$ $87.2$ $1000$
MGPS [moufad2024variational] $20.5$ $7.7$ $9.7$ $13.9$ $10.5$ $24.3$ $11.7$ $8.2$ $29.0$ $1764$
@@
Palette [saharia2022palette] $21.0$ $4.7$ $5.8$ $17.0$ $5.3$ $5.5$ $12.5$ 6.6 $25.5$ $63$
InDI [delbracio2023inversion] $21.4$ $21.6$ $22.8$ $17.7$ $22.6$ $25.0$ $13.7$ $35.5$ $58.7$ $50$
InvFusion $22.7$ 4.2 4.2 $17.1$ 5.0 5.0 $15.1$ $7.1$ 7.1 $63$
MMSE predictor (Training-based)
Palette [saharia2022palette] $23.5$ - - $19.5$ - - $15.5$ - - 11
InvFusion 25.0 - - 19.8 - - 17.9 - - 11
@@
DDRM [kawar2022denoising] $21.2$ $26.1$ $26.2$ $12.2$ $30.5$ $30.5$ $10.6$ $32.9$ $32.9$ $25$
DDNM [wang2022zero] $22.5$ $69.5$ $69.4$ $15.1$ $95.3$ $95.3$ $11.3$ $115.3$ $115.3$ $100$
DPS [chung2023diffusion] $18.3$ $9.0$ $12.9$ $10.4$ $15.2$ $17.7$ $9.5$ $12.8$ $12.8$ 10001000*
[…]
DAPS [zhang2024improving] $17.3$ $125.7$ $62.8$ $15.4$ $16.3$ $19.7$ $11.7$ $28.3$ $37.0$ $1000$
MGPS [moufad2024variational] $19.4$ $14.2$ $18.0$ $13.3$ $18.5$ $25.0$ $11.7$ $13.8$ $18.9$ $1764$
@@
Palette [saharia2022palette] $19.6$ $6.6$ $7.0$ $15.6$ $6.2$ $6.2$ $12.3$ $6.7$ $11.4$ $63$
InDI [delbracio2023inversion] $20.6$ $23.3$ $24.6$ $17.1$ $21.0$ $21.0$ $14.5$ $30.5$ $40.0$ $50$
InvFusion $20.9$ 5.6 5.6 $15.6$ 5.9 5.9 $14.5$ 6.3 6.3 $63$
MMSE predictor (Training-based)
Palette [saharia2022palette] $22.3$ - - $18.3$ - - $15.3$ - - 11
InvFusion 23.4 - - 18.4 - - 17.3 - - 11
    Methods that use the network derivative, using more computation and memory per NFE.
@@ 4 Experiments
We evaluate the advantage of InvFusion over existing inverse problem solvers on $64\times 64$ images from the FFHQ [ffhq] and ImageNet [imagenet] datasets. In these experiments, we conduct a comparative analysis of InvFusion against alternative methods that can operate on several types of degradations with a single model. This includes zero-shot methods, utilizing an comparable unconditional model. Additionally, we evaluate our approach against a “Palette”-style methodology [saharia2022palette] that incorporates the ${\bm{H}}^{\dagger}{\mathbf{y}}$ as input without additional degradation information. We train these models ourselves using the same underlying architecture, to enable a comprehensive assessment of our method’s capabilities disentangled from architectural considerations. The models trained on ImageNet are also class-conditional, and use CFG [ho2022classifier] with a factor of $2.0$ for sampling.
@@
We train and evaluate our model on several categories of inverse problems; strided motion blur (combining various degrees of super-resolution and de-blurring), in-painting by leaving only patches of various sizes, and a general matrix degradation operator. These degradations each have computationally efficient implementations for applying both $\bm{H}$} and $\bm{H}$$^{\dagger}$. The parameters of each degradation are randomized, to create a large corpus of possible degradations (as detailed in App.˜A). All conditional models are also trained on unconditional generation by using a degradation which always outputs zeros. Figure 3 shows qualitative comparisons of different inverse problem solving methods for the two datasets. All training-based approaches use the deterministic sampler from EDM [karras2022elucidating] along with identical seeds, highlighting the effects of different models.
@@
For quantitative comparisons, we use PSNR to measure the image fidelity, i.e. how close the restored output is to the original. To measure image realism, we use 10K FID [fid]. These two metrics are at odds according to the Perception-Distortion-Tradeoff [blau2018perception], and different methods excel at either of the metrics. We find that some restoration algorithms seem to sacrifice consistency with the measurements $\mathbf{y}$} to accommodate better image realism, in effect ignoring the inverse problem. To measure the image realism of valid solutions to the given inverse problem, we offer to measure the FID of generated samples that have been projected to be consistent with the measurements $\mathbf{y}$}, which we refer to as CFID. In this way, we penalize models for the distance between the generated sample and the nearest valid solution to the formulated problem. We also include LPIPS [lpips] measures in App.˜D. Tables 1 and 2 show our qualitative analysis on the FFHQ [ffhq] and ImageNet [imagenet] datasets respectively. Our model achieves the best CFID among all training-based and zero-shot methods, along with the best FID in all but a single case, suggesting it is SOTA in generating samples from the posterior for our setting. This superior quality is attained despite using far less computations than nearly all zero-shot methods, as quantified by Neural Function Evaluations (NFEs). Our InvFusion MMSE estimator based on the same model achieves the best PSNR of all models, while the InvFusion posterior sampler is also among the highest in PSNR.
@@
Interestingly, the gap in performance between InvFusion and Palette highly varies between different degradation families. In problems like in-painting, the degradation is easily inferrable by the network, as the ${\bm{H}}^{\dagger}{\mathbf{y}}$ input to the network contains masked regions (see Fig.˜3). Thus in in-painting, we see a small advantage for InvFusion over Palette. On the other hand, in motion-blur and even more so on general matrix degradations, Palette struggles to infer the exact degradation, leading to a drop in both PSNR and CFID, sometimes, even lower than some zero-shot methods which are degradation-aware. This supports our hypothesis that knowledge of the degradation operator, whether inferred or explicit, is critical for correct restoration.
@@
Previous comparisons focused on smaller images ($64\times 64$) to enable efficient evaluations of many methods and degradations, some of which are computationally expensive. To demonstrate InvFusion’s scalability and performance at higher resolutions, we conducted additional experiments on strided motion blur and patch in-painting using $256\times 256$ images from FFHQ. The complete results are presented in App.˜D, with representative examples shown in Fig.˜1.
@@
Table 3: Comparison of restoration on a degradation that did not appear in training. InvFusion demonstrates strong adaptation capabilities through its degradation-aware architecture.
@@
DDRM [kawar2022denoising] $11.6$ $24.9$ $24.9$
DDNM [wang2022zero] $12.7$ $75.8$ $75.8$
DPS [chung2023diffusion] $10.1$ $8.8$ $14.1$
[…]
DAPS [zhang2024improving] $12.1$ $40.8$ $39.5$
@@
Palette [saharia2022palette] $11.0$ $19.9$ $19.9$
InvFusion $11.9$ 5.9 5.9
MMSE predictor (Training-based)
Palette [saharia2022palette] $13.6$ - -
InvFusion 14.7 - -
Table 4: Unconditional sampling FID.
@@
Baseline [crowson2024scalable] 6.66 8.98
Palette [saharia2022palette] 6.56 8.26
InvFusion 6.36 8.17
Figure 4: Using InvFusion NPPC. By being degradation-aware, the model can be trained to predict the MMSE along with several leading principal components ${\mathbf{w}}_{i}$ (left, contrast enhanced) for many degradations.
@@ 4.1 Out-Of-Distribution Degradations and Unconditional Sampling
To measure the generalization capabilities of different methods we evaluate performance on a degradation that is slightly outside of the training distribution. In this case, we perform out-painting from a single rectangular area in the image. The results in Tab.˜3, show that Palette completely fails to generalize, despite the similarity to the in-painting task the model has been trained on and the ease at which the degradation operator can be inferred from the input. On the other hand, InvFusion does quite well on this problem, outperforming all zero-shot methods, which are not penalized by the out-of-distribution (OOD) degradation, as they are not trained on any degradation. Notably, the matrix degradations demonstrated in Tabs. 1 and 2 underscore the generalization capabilities of InvFusion, as the range of possible degradations is so huge that it is near certain that none of the degradations used in test-time have appeared in training. Finally, we also evaluate all trained models for unconditional generation (on which they had been trained, by using ${\bm{H}}\in R^{0\times D}$). Surprisingly, we find that both conditional models outperform the unconditional baseline, with InvFusion leading with approximately $5\%$ improvement over the baseline (unconditional diffusion). This is despite the conditional models having no additional capacity in unconditional generation, as both the additional model inputs and the Feature Degradation Layer are always output zeros in this setting. We hypothesize that being exposed to partial image information during training helps the diffusion model converge. This indicates that a single InvFusion model can be well adapted for both restoration and unconditional generation.
@@ 4.2 Noisy Degradations
Following the noiseless linear case explored in earlier sections, we also conduct experiments on noisy inverse problems. We retrain the FFHQ64 model, incorporating additive white Gaussian noise with a standard deviation of $\sigma_{n}=0.1$ in the measurements $\mathbf{y}$}. The InvFusion architecture remains unchanged, using only the transforms $\bm{H}$} and $\bm{H}$$^{\dagger}$ of the degradation – as adding randomly sampled noise at each layer could lead to accumulation of noise. Projection operators are not well-defined for noisy degradations, so only PSNR and FID are evaluated for the noisy case. Our findings in Tab.˜11 (App.˜D) demonstrate the advantage of InvFusion over alternative methods in the noisy settings.
@@ 5 Neural Posterior Principal Components
Beyond the improvements in restoration fidelity and realism, using an operator-aware architecture unlocks several capabilities that are useful for down stream applications. A notable example, is the computation of Neural Posterior Principal Components (NPPCs) [nehme2024uncertainty], which are efficient estimations of the posterior distribution’s largest principal vectors. These direction are meaningful for uncertainty quantification among many other tasks [manor2023posterior, elata2025psc], and may have high implications for medical analysis [belhasin2023principal, elata2025adaptive]. In the original work by nehme2024uncertainty, a network is trained to approximate the posterior principal components for a single pre-defined degradation. Using InvFusion, we can use the same approach to scale up NPPC training to multiple degradations at once, enabling flexible adaptation to the degradation at runtime. We experiment with applying the exact same NPPC loss from [nehme2024uncertainty], using the InvFusion architecture to condition the network on the degradation operator $\bm{H}$} for the in-painting of a single area of the image. The examples in Fig.˜4 shows that InvFusion meaningfully extends NPPC computation for multiple degradation, making a single model viable for a wide range of problems.
@@ 6 Discussion, Limitations and Conclusion
Despite achieving state-of-the-art performance in posterior sampling and MMSE estimation, InvFusion faces several important limitations. A primary constraint lies in the model’s training scope – although InvFusion demonstrates remarkable adaptability to various degradations, it can only be trained on a finite set of degradation types, beyond which the model may under-perform. Even in the linear case, the space of possible degradations scales quadratically with image dimensions, creating fundamental constraints that would be impractical to address solely through increased model capacity and training time. Additionally, while our current results focus exclusively on linear degradations, we anticipate that InvFusion could potentially generalize to non-linear degradations when appropriately trained, though this remains to be empirically validated. This in turn would enable the use of InvFusion for latent diffusion, in which non-linear operators apply or approximate the encoder and decoder. A final consideration is the computational and memory intensity of repeated degradation applications, particularly during the training phase. Although this computational overhead may be relatively minor compared to certain zero-shot methods that require additional steps or model derivatives, optimizing these resource requirements represents a key area for future improvement.
This work introduces InvFusion, a novel architectural framework that bridges the long-standing gap between training-based and zero-shot approaches in solving inverse problems with diffusion models. By incorporating the degradation operator directly into the network architecture through the attention mechanism, our method achieves state-of-the-art performance while maintaining the flexibility to handle diverse degradation scenarios. The empirical results demonstrate superior performance on multiple datasets and across various inverse problems, while offering computational efficiency comparable to many existing methods. Beyond its primary application, InvFusion’s capability as a general MMSE estimator and its potential for NPPC estimation opens new avenues for downstream applications. While certain limitations remain, this work represents a significant step forward in the field of accurate and efficient inverse problem solving.
Acknowledgments and Disclosure of Funding
@@
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)), and by the National Research Foundation of Korea under Grant RS-2024-00336454.
Appendices
Table 5: Architecture Hyperparameters
Experiment Patch Depths Widths Attention Type
FFHQ64 $(2,2)$ $[2,2,8]$ $[384,768,1536]$ $\text{NAttn}7,\text{NAttn}7,\text{GAttn}]$
ImageNet64 $(2,2)$ $[2,2,8]$ $[384,768,1536]$ $\text{NAttn}7,\text{NAttn}7,\text{GAttn}]$
FFHQ256 $(4,4)$ $[2,2,4,2]$ $[192,384,768,1536]$ $\text{NAttn}5,\text{NAttn}7,\text{GAttn},\text{GAttn}]$
Table 6: Training Hyperparameters
Experiment Iters Batch Aug Prob Label Dropout Mixed Precision
FFHQ64 $2^{16}$ $512$ $0.12$ - fp16
ImageNet64 $2^{18}$ $512$ $0.0$ $0.1$ fp16
FFHQ256 $2^{18}$ $256$ $0.12$ - bf16